# MaxInfo: A Training-Free Key-Frame Selection Method Using Maximum Volume for Enhanced Video Understanding

### v2

@@ Abstract
Modern Video Large Language Models (VLLMs) often rely on uniform frame sampling for video understanding, but this approach frequently fails to capture critical information due to frame redundancy and variations in video content. We propose MaxInfo, a training-free method based on the maximum volume principle, which selects and retains the most representative frames from the input video. By maximizing the geometric volume formed by selected embeddings, MaxInfo ensures that the chosen frames cover the most informative regions of the embedding space, effectively reducing redundancy while preserving diversity. This method enhances the quality of input representations and improves long video comprehension performance across benchmarks. For instance, MaxInfo achieves a 3.28% improvement on LongVideoBench and a 6.4% improvement on EgoSchema for LLaVA-Video-7B. It also achieves a 3.47% improvement for LLaVA-Video-72B. The approach is simple to implement and works with existing VLLMs without the need for additional training, making it a practical and effective alternative to traditional uniform sampling methods.
Training-Free, Plug-and-Play, VLLMs, Video Understanding
Figure 1: Reasons for not being able to answer the correct answer at long video benchmark (Video-MME).
@@ 1 Introduction
Large language models (LLMs) such as GPT (Floridi & Chiriatti, 2020; Achiam et al., 2023), LLaMA (Touvron et al., 2023; Dubey et al., 2024), Qwen (Bai et al., 2023; Yang et al., 2024), and Mistral (Jiang et al., 2023) have revolutionized tasks like text generation, summarization, and reasoning. Recent advancements in multimodal large language models (MLLMs) (Liu et al., 2024; Goncharova et al., 2024) have extended these capabilities to include processing images, videos, and audio, enabling responses across diverse modalities. Video understanding, in particular, has garnered significant attention due to its complex, multi-dimensional nature and broad range of applications.
While models like LLaVA-Video (Zhang et al., 2024c), VideoLLaMA 2 (Cheng et al., 2024), MiniCPM-V 2.6 (Yao et al., 2024), and InternVL (Chen et al., 2024) have made strides in video understanding, they struggle with long videos due to the diversity and redundancy of video content. Uniform frame sampling, a widely used approach, often fails to capture the most informative frames, leading to missing critical details and model performance decrease. As in Figure 1 illustrate this challenge using the Video-MME (Fu et al., 2024) benchmark. As shown, uniform sampling can overlook key information essential to understanding the video, as frames critical to the answer may not be selected.
Existing approaches attempt to address these challenges by either increasing input sequence length or compressing video information. Models like LongVU (Shen et al., 2024) and MovieChat (Song et al., 2024) compress tokens per frame, while others, such as Qwen2-VL (Wang et al., 2024a) and Gemini 1.5 Pro (Team et al., 2024), process longer sequences, supporting up to 32K and 10 million tokens, respectively. However, these solutions either incur substantial computational overhead or risk losing critical temporal information. Balancing the trade-off between efficiency and accuracy remains a key challenge in long video understanding.
@@
Experimental results demonstrate that MaxInfo significantly improves the performance of existing models, such as LLaVA-Video (Zhang et al., 2024c), InternVL2 (Chen et al., 2024) and Qwen2-VL (Wang et al., 2024a), achieving better results on benchmarks like Video-MME (Fu et al., 2024) and EgoSchema (Mangalam et al., 2023). By overcoming the limitations of uniform sampling, MaxInfo enables more accurate and robust video understanding, ensuring that critical information from long videos is effectively captured and utilized in multimodal applications.
@@ 2 Related Works
Video Large Language Models (VLLMs). Video understanding has become a focal area of research, with numerous models excelling at video comprehension tasks. These tasks typically involve converting videos into image frames and inputting them into VLLMs. Existing approaches fall into two main categories:
@@
Using query-based models like Q-Former (Li et al., 2023) to extract critical visual features from image frames, which are then processed by VLLMs (Li et al., 2024; Xenos et al., 2024). Encoding frame sequences with models such as CLIP (Radford et al., 2021), DINO (Caron et al., 2021), and Siglip (Zhai et al., 2023), and feeding the resulting embeddings into VLLMs (Shen et al., 2024; Jin et al., 2024a, b; Wang et al., 2024a; Chen et al., 2024; Goncharova et al., 2024). While these methods emphasize visual feature extraction and text-image semantic understanding, they often rely on uniform frame sampling or similarity-based techniques, which can overlook critical information, especially in long videos with diverse content.
Long Video Understanding. Understanding long videos poses significant challenges, primarily due to the need to balance computational efficiency with preserving critical temporal and contextual information. To address this, various strategies have been proposed:
@@
Reducing sequence length: Methods like Video-LaVIT (Jin et al., 2024b) and LongVU (Shen et al., 2024) use cosine similarity or clustering to filter redundant frames, while MovieChat (Song et al., 2024) applies similarity thresholds for frame selection. Token compression: SlowFast-LLaVA (Xu et al., 2024) compresses visual tokens, and Chat-UniVi (Jin et al., 2024a) extracts key event tokens to reduce redundancy. Extended input lengths: Models such as Qwen2-VL (Wang et al., 2024a) and Gemini 1.5 Pro (Team et al., 2024) handle extended token lengths to process long videos, albeit with high computational costs. Despite these advancements, current methods often depend on arbitrary thresholds, fixed compression schemes, or uniform sampling, which may fail to capture the diverse and critical content of long videos effectively.
@@
Information Maximization Techniques. Information maximization has been widely used in feature selection and dimensionality reduction across various domains. Methods like mRMR (Peng et al., 2005) and MMD (Vasconcelos, 2002) improve model performance by selecting features with high relevance and low redundancy, while MOI (Sindhwani et al., 2004) enhances classification by focusing on the most useful feature subsets. Among these, the maximum volume (MaxVol) algorithm (Goreinov et al., 2010; Mikhalev & Oseledets, 2018; Sozykin et al., 2022) stands out for its ability to maximize information by selecting linearly independent rows of a matrix, ensuring the chosen subset spans the most informative subspace. However, these methods are generally designed for static datasets and are not directly applicable to dynamic video frame selection.
@@
In this paper, we extend the principles of MaxVol to video understanding, introducing a keyframe extraction framework tailored for VLLMs. Unlike existing approaches, our method dynamically selects diverse and representative frames. By maximizing the geometric volume of frame embeddings, MaxInfo ensures the selected frames are both diverse and highly informative. Furthermore, we advance this framework with a scene-aware algorithm, enabling finer-grained selection of critical frames for each scene. These contributions address the limitations of uniform sampling, providing a principled and effective training-free solution for long video understanding.
@@ 3.1 Overview
Let the sampled frames be $\mathbf{I}\;=\;\{\,i_{1},\;i_{2},\;\ldots,\;i_{n}\}.$ We extract each frame’s visual representation via a CLIP-based ViT (Radford et al., 2021) and retain only the [CLS] token. Stacking these tokens yields
$\mathbf{Q}\;=\;\begin{bmatrix}q_{1}\\[4.0pt] q_{2}\\[4.0pt] \vdots\\[2.0pt] q_{n}\end{bmatrix}\;\in\;\mathbb{R}^{n\times d},$ (1)
@@ 3.3 Rectangular MaxVol Frame Selection
On the next step we identify the “most informative” subset of rows of $\mathbf{Q}_{s}$, i.e., a set of frames which corresponding rows span the overall distribution of frames. To do this, we use the rectangular MaxVol algorithm (Mikhalev & Oseledets, 2018) to evaluate a submatrix of maximal volume in $\mathbf{Q}_{s}$.
@@
[…]
@@
These indices $\mathbf{r}$ then specify the frames $i_{n}$ we deem most representative.
@@
$\mathbf{S}\;=\;\mathbf{Q}_{s}(\mathbf{r},:)$ (6)
is an $r\times s$ matrix containing the “diverse, high-information” frames. We feed only these $r$ frames into the downstream VLLM:
@@
Since $r\ll n$ in most cases, this yields a more efficient inference process without sacrificing essential visual context. In practice, VLLMs typically process a fixed number of frames k (e.g., every 1, 32, 64, 128 frame). Instead of directly selecting k frames or reducing the number of frames the network processes, we propose two versions of MaxInfo implementation: fast and slow. In the fast version, we apply MaxVol to the default number of frames and reduce it (e.g., reducing 256 frames to 180). In the slow version, we apply MaxVol to a larger set of frames (e.g., starting with 1024 frames and reducing it to 256), ensuring a richer pool of information for frame selection. We then apply CLIP embeddings and the MaxVol algorithm to intelligently refine this selection, maximizing the diversity and informativeness of the final k frames passed to the network. This ensures that VLLMs receive the most representative frames, improving their ability to understand long-form video content efficiently.
@@ 3.4 Fast and Slow Version
3.5 Scene-Aware MaxInfo
Videos often contain multiple scenes with similar visual content per scene, making Chunk-Based MaxInfo selection suboptimal. If applied globally, MaxInfo may discard crucial frames due to artificial embedding similarities across different scenes.
To address this, we introduce Scene-Aware MaxInfo, a very simple and effective modification. Instead of performing complex scene detection, we uniformly split the video into 32 equal-sized chunks and apply MaxVol independently within each chunk. This ensures that each segment is well-represented while maintaining computational simplicity.
Chunk-Based MaxVol Selection
Given $n$ uniformly sampled frames, we split them into 32 contiguous chunks:
$\mathbf{I}=\bigcup_{i=1}^{32}\mathbf{I}^{(i)},\quad\mathbf{I}^{(i)}=\{i_{j}\mid j\in\text{chunk }i\}.$ (9)
For each chunk, we extract CLIP embeddings, apply SVD for dimension reduction, and run MaxVol to select the most representative frames.
@@ 3.6 Summary
Our pipeline can be summarized with an Algorithm 1. Because our method is training-free and focuses purely on data selection, it can be plugged into diverse video-focused VLLMs to reduce cost and improve performance on long-video tasks.
@@
$q_{n}=\mathrm{flatten}\bigl{(}\mathrm{clip\_model}(i_{n})\bigr{)},\mathbf{Q}=\begin{bmatrix}q_{1}\\ q_{2}\\ \vdots\\ q_{n}\end{bmatrix}\in\mathbb{R}^{n\times d}.$
@@
[…]
@@ 4 Experiments
4.1 Implementation Details
In order to evaluate the contribution of MaxInfo to video understanding, we employed widely-used video understanding benchmarks, covering short-video tasks (such as MVBench (Li et al., 2024)) and medium-to-long video tasks (such as EgoSchema (Mangalam et al., 2023), Video-MME (Fu et al., 2024), and LongVideoBench (Shen et al., 2024)). For complete fairness, we only compared improved versions of the model against itself without MaxVol with freezing generation parameters, seed and prompt. We mostly focus on longer videos because better frame selection plays a bigger role in longer, more complex videos, whereas shorter ones intuitively work well with uniform sampling due to their lower information content and complexity.
For the general multiple-choice question-answering evaluation, we follow the official guidelines to construct the instructions using the provided questions and options. We added a prompt to the question and options like ”Respond with only the letter (A, B, C, or D) of the correct option.” for LongVideoBench (Wu et al., 2024), Video-MME (Fu et al., 2024), MLVU (Zhou et al., 2024) and MVBench (Li et al., 2024) or ”Answer with the option’s letter from the given choices directly and only give the best option.” for EgoSchema (Mangalam et al., 2023). We follow the original benchmarks setup to calculate the final scores, and we also align our evaluation protocols with other evaluation toolkits, such as lmms-eval (Zhang et al., 2024a).
@@ 4.2 Main Results
Overall Performance. Table 1, Table 2 and Table 3 present the performance gains achieved by integrating MaxInfo into existing InternVL (Chen et al., 2024), Qwen2-VL (Wang et al., 2024a) and LlaVA-Video (Zhang et al., 2024c) models. The experimental results show that MaxInfo Block exhibits significant performance improvement on a wide range of models. In particular, in the LLaVA-Video-7B and Qwen-VL-2B models, the introduction of MaxInfo Block improves the accuracy by 0.9%/1.7%, 6.4%, 3.3% and 1.4%/1.2%, 2.3%, 1.5% in VideoMME, EgoSchema, and LongVideoBench, respectively, which is significantly better than the versions without the block. This improvement not only validates the effectiveness of MaxInfo Block, but also provides new ideas for future research on video understanding tasks.
Table 1: Comparing Video-LLMs with and without MaxInfo as an additional module. Except for ours (-MaxInfo) and the models with *, all results are copied from the related papers of benchmarks or models. Base models uses uniform frames sampling. For the Video-MME benchmark, we report results under two standard settings: without subtitles (wo. sub.) and with subtitles (w. sub.). The % accuracy sign is omitted for clarity.
@@
LLaVA-Video (Zhang et al., 2024c) 7B 63.3/69.7(64) 57.3(64) 58.2(64)
+ MaxInfo 7B 64.2/71.464→(6,64) 63.7128→(64,64) 61.5128→(1,64)
$\triangle$ +0.9%/+1.7% +6.4% +3.3%
LLaVA- (Zhang et al., 2024c) 72B 70.5/76.9(64) 65.6(64) 61.9(64)
+ MaxInfo 72B 70.2/77.664→(6,64) 69.4128→(64,64) 64.9128→(64,1)
$\triangle$ -0.3%/+0.7% +3.8% +3%
Qwen2-VL (Wang et al., 2024a) 2B 55.6/60.4(786) 54.9(180) 47.3(256)
+ MaxInfo 2B 57.0/61.6256→(254,4) 57.2180→(180,12) 48.8256→(224,1)
$\triangle$ +1.4%/+1.2% +2.3% +1.5%
Qwen2-VL (Wang et al., 2024a) 7B 63.3/69.0(768) 66.7(180) 53.7(256)
+ MaxInfo 7B 62.1/70.0256→(254,4) 64.3180→(180,12) 55.7256→(224,1)
$\triangle$ -1.2/+1.0% +2.4% +2.0%
Table 2: Comparison of the InternVL2 model series with and without MaxInfo as an additional module. InternVL2 uses uniform frame sampling and has a limited context size. Therefore, the maximum number of frames for evaluation across all benchmarks was limited to 32. The maximum number of input and output frames for the MaxInfo module is set to 128 and 32 respectively. For the LongVideoBench benchmark, we report results under a default setting for video LLMs, i.e. without subtitles. The % accuracy sign is omitted for clarity.
Model Size MVBench Egoschema LongVideoBench
InternVL2 (Chen et al., 2024) 1B 55.42 34.02 43.38
+ MaxInfo(128,32) 1B 56.07 34.11 42.78
$\triangle$ +0.65% +0.09% -0.6%
InternVL2 (Chen et al., 2024) 2B 59.72 46.10 43.97
+ MaxInfo(128,32) 2B 59.02 46.33 47.27
$\triangle$ -0.7% +0.23% +3.3%
Table 3: Performance comparison between MaxInfo Block and other key frame selection methods on the Video-MME-Long benchmark. Results are measured without adding subtitles. $Q$ refers to the Qwen2-VL-7B model and $L$ refers to the LLaVA-Video-7B model. The % accuracy sign is omitted for clarity.
Model Size frames Acc.
Video-Tree (Wang et al., 2024c) - - 54.2/-
Long-VU (Shen et al., 2024) 7B 1fps -/60.6
Chat-Univi (Jin et al., 2024a) 7B 64 35.8/-
Q+MaxInfo 7B 256 $\rightarrow$ (254, 4) 51.8/63.9
L+MaxInfo 7B 64 $\rightarrow$ (64, 6) 54.6/65.9
MLVU Benchmark. Table 4 evaluates both MaxInfo and the benefits of Chunk-Based Scene-Awareness on the MLVU benchmark, demonstrating consistent improvements over uniform sampling. While these improvements are modest, it is important to highlight that the method is exceptionally simple and entirely training-free. This simplicity underscores the potential of even basic scene-awareness to enhance performance, suggesting that more sophisticated scene detection techniques could lead to further gains. Such results indicate that scene-awareness is a promising direction for improving video understanding tasks.
Table 4: Comparison of performance on the MLVU benchmark between the original versions of the models in the Qwen-VL family and their variants with the MaxInfo and Naive Scene-Aware (NSA) additional modules. The original models use uniform sampling with a maximum of 128 frames. The additional modules use from 128 to 1024 frames per input. The best results are highlighed in bold, second best are underlined.
@@
4.3 Performance Analysis: MaxInfo vs. Uniform Sampling
To better understand the strengths and trade-offs of MaxInfo, we analyzed per-task accuracy across multiple benchmarks.
Figure 3: Accuracy comparison between Uniform Sampling and MaxInfo on Video-MME (Qwen2-VL-2B).
Figure 4: Accuracy comparison between Uniform Sampling and MaxInfo on LongVideoBench (Qwen2-VL-2B).
Figure 5: Accuracy comparison between Uniform Sampling and MaxInfo on MLVU (Qwen2-VL-7B) with 64 final and 1024 initial frames.
Our results, as shown in Figure 3, Figure 4, and Figure 5, indicate that MaxInfo outperforms uniform sampling on tasks requiring high information density (e.g., counting, summarization, and spatial reasoning). However, for tasks requiring fine-grained temporal reasoning (e.g., object recognition and action perception), uniform sampling retains a slight advantage. These findings highlight a fundamental trade-off between information maximization and temporal continuity, suggesting potential directions for hybrid selection methods.

### v3

@@ Abstract
Modern Video Large Language Models (VLLMs) often rely on uniform frame sampling for video understanding, but this approach frequently fails to capture critical information due to frame redundancy and variations in video content. We propose MaxInfo, the first training-free method based on the maximum volume principle, which is available in Fast and Slow versions and a Chunk-based version that selects and retains the most representative frames from a video. By maximizing the geometric volume formed by selected embeddings, MaxInfo ensures that the chosen frames cover the most informative regions of the embedding space, effectively reducing redundancy while preserving diversity. This method enhances the quality of input representations and improves long video comprehension performance across benchmarks. For instance, MaxInfo achieves a 3.28% improvement on LongVideoBench and a 6.4% improvement on EgoSchema for LLaVA-Video-7B. Moreover, MaxInfo boosts LongVideoBench performance by 3.47% on LLaVA-Video-72B and 3.44% on MiniCPM4.5. The approach is simple to implement and works with existing VLLMs without the need for additional training and very lower latency, making it a practical and effective alternative to traditional uniform sampling methods. Our code are available at https://github.com/FusionBrainLab/MaxInfo.git
Figure 1: Reasons why the Uniform Sampling approach cannot answer the correct answer in long videos. An example of MaxInfo’s sampling approach.
@@ 1 Introduction
Large language models (LLMs) such as GPT [gpt-3, gpt4], LLaMA [llama2, llama3], Qwen [qwen, qwen2], and Mistral [jiang2023mistral] have revolutionized tasks like text generation, summarization, and reasoning. Recent advancements in multimodal large language models (MLLMs) [llava, omnifusion] have extended these capabilities to include processing images, videos, and audio, enabling responses across diverse modalities. Video understanding, in particular, has garnered significant attention due to its complex, multi-dimensional nature and broad range of applications.
While models like LLaVA-Video [llava-video], VideoLLaMA 2 [cheng2024videollama2], MiniCPM-V 2.6 [yao2024minicpmv], and InternVL [internvl] have made strides in video understanding, they struggle with long videos due to the diversity and redundancy of video content. Uniform frame sampling, a widely used approach, often fails to capture the most informative frames, leading to missing critical details and model performance decrease. As shown in Figure 1 illustrate this challenge using the Video-MME [video-mme] benchmark. As shown, uniform sampling can overlook key information essential to understanding the video, as frames critical to the answer may not be selected.
Existing approaches attempt to address these challenges by either increasing input sequence length or compressing video information. Models like LongVU [longvu] and MovieChat [song2024moviechat] compress tokens per frame, while others, such as Qwen2-VL [qwen2-vl] and Gemini 1.5 Pro [team2024gemini], process longer sequences, supporting up to 32K and 10 million tokens, respectively. However, these solutions either incur substantial computational overhead or risk losing critical temporal information. Balancing the trade-off between efficiency and accuracy remains a key challenge in long video understanding.
@@ 2 Related Works
Video Large Language Models (VLLMs). Video understanding has become a focal area of research, with numerous models excelling at video comprehension tasks. These tasks typically involve converting videos into image frames and inputting them into VLLMs. Existing approaches fall into two main categories: Using query-based models like Q-Former [blip2] to extract critical visual features from image frames, which are then processed by VLLMs [mvbench, xenos2024vllms]. Encoding frame sequences with models such as CLIP [clip], DINO [dino], and Siglip [Siglip], and feeding the resulting embeddings into VLLMs [longvu, Chat-univi, video-lavit, qwen2-vl, internvl, omnifusion, mvbench, chen2024sharegpt4video]. While these methods emphasize visual feature extraction and text-image semantic understanding, they often rely on uniform frame sampling or similarity-based techniques, which can overlook critical information, especially in long videos with diverse content.
@@
Long Video Understanding. Understanding long videos poses significant challenges, primarily due to the need to balance computational efficiency with preserving critical temporal and contextual information. To address this, various strategies have been proposed, Reducing sequence length: Methods like Video-LaVIT [video-lavit] and LongVU [longvu] use cosine similarity or clustering to filter redundant frames, while MovieChat [song2024moviechat] applies similarity thresholds for frame selection. Token compression: SlowFast-LLaVA [slowfast] compresses visual tokens, and Chat-UniVi [Chat-univi] extracts key event tokens to reduce redundancy. Extended input lengths: Models such as Qwen2-VL [qwen2-vl] and Gemini 1.5 Pro [team2024gemini] handle extended token lengths to process long videos, albeit with high computational costs.
In addition, numerous keyframe extraction algorithms have been proposed, such as LongVA [zhang2024long], Frame-Voyager [yu2024frame], AKS [tang2025adaptive], VideoTree [videotree], M-LLM [hu2025m], BOLT [liu2025bolt], VSLS [guo2025logic], AdaReTaKe [wang2025adaretake], Q-Frame [zhang2025q], GenS [yao2025generative], and ViLaMP [cheng2025scaling], all of which have demonstrated remarkable performance across various long video understanding tasks.
Despite these advancements, current methods often depend on arbitrary thresholds, fixed compression schemes, or uniform sampling, which may fail to capture the diverse and critical content of long videos effectively.
@@
Information Maximization Techniques. Information maximization is widely used for feature selection and dimensionality reduction. Methods such as mRMR [peng2005feature] and MMD [MMD] improve model performance by selecting features with high relevance and low redundancy, while MOI [MOI] focuses on the most informative feature subsets to enhance classification. The maximum volume (MaxVol) algorithm [goreinov2010find, Rectmax, sozykin2022ttopt] selects linearly independent rows of a matrix to cover the most informative subspace.
@@
In this paper, we extends the MaxVol principle to video understanding and proposes a keyframe extraction framework tailored for VLLMs. Unlike existing methods, our approach dynamically selects diverse and representative frames. By maximizing the geometric volume of frame embeddings, MaxInfo ensures that the selected frames are both informative and diverse, and, combined with a scene-aware algorithm, enables fine-grained selection of keyframes for each video segment, providing an efficient, training-free solution for long video understanding.
@@ 3.1 Overview
Let the sampled frames be $\mathbf{I}\;=\;\{\,i_{1},\;i_{2},\;\ldots,\;i_{n}\}.$ We extract each frame’s visual representation via a CLIP-based ViT [clip] and retain only the [CLS] token. Stacking these tokens yields
$\mathbf{Q}=\begin{bmatrix}q_{1}&q_{2}&\cdots&q_{n}\end{bmatrix}^{\top}$ (1)
@@ 3.3 Rectangular MaxVol Frame Selection
On the next step we identify the “most informative” subset of rows of $\mathbf{Q}_{s}$, i.e., a set of frames which corresponding rows span the overall distribution of frames. To do this, we use the rectangular MaxVol algorithm [Rectmax] to evaluate a submatrix of maximal volume in $\mathbf{Q}_{s}$.
@@
$\text{rect-vol}(\mathbf{A})\;=\;\sqrt{\det(\mathbf{A}\,\mathbf{A}^{T})}$ (4)
Maximizing $\text{rect-vol}(\mathbf{A})$ with respect to the selection of rows corresponds to identifying the subset of frames that best preserves the variation in the data. We denote the selected row indices as
@@
where indices $\mathbf{r}$ then specify the frames $i_{n}$ we deem most representative.
@@
$\mathbf{S}\;=\;\mathbf{Q}(\mathbf{r},:)$ (6)
is an $r\times n$ matrix containing the diverse, high-information frames. We feed only these $r$ frames into the downstream VLLM:
@@
Given that the value of r is much smaller than n (i.e., $r\ll n$) in most application scenarios, the design significantly improves the inference efficiency while effectively avoiding the loss of critical visual context information. We innovatively propose the MaxInfo algorithm framework scene-aware MaxInfo and fast and slow versions of the implementation scheme. Experimental results show that both exhibit significant enhancements.
@@ 3.4 Fast and Slow Version
3.5 Chunk-Based MaxInfo
Videos often contain multiple scenes with visually similar frames, making global frame selection suboptimal. Applied across the entire video, MaxInfo may also discard important frames due to spurious embedding similarities between different scenes.
To address this, we propose Chunk-Based MaxInfo, a simple yet effective modification. We uniformly divide the video into $M$ equal-sized chunks and apply MaxInfo independently within each chunk. This ensures that every segment is adequately represented while keeping the procedure computationally efficient.
Formally, given $n$ uniformly sampled frames, we split them into $M$ contiguous chunks:
$\mathbf{I}=\bigcup_{i=1}^{M}\mathbf{I}^{(i)},\quad\mathbf{I}^{(i)}=\{i_{j}\mid j\in\text{chunk }i\}.$ (9)
For each chunk, we extract CLIP embeddings, apply SVD for dimensionality reduction, and run MaxVol to select the most representative frames. In our experiments, we set $M=32$ for simplicity, though any choice of $M$ is possible and can be tuned to balance representation quality and computational cost.
@@ 3.6 Summary
Our pipeline can be summarized with an Algorithm 1. It is a training-free and plug-and-play algorithm that can be integrated into any model of VLLMs.
@@
$q_{n}=\mathrm{flatten}\bigl(\mathrm{clip}(i_{n})\bigr),\mathbf{Q}=\begin{bmatrix}q_{1}&q_{2}&\cdots&q_{n}\end{bmatrix}^{\top}$
@@
4:MaxVol Selection: Run $\text{rect\_maxvol}(\mathbf{Q_{s}},\mathrm{tol})$ to find pivot indices:
@@ 4 Experiments
Overall. In order to evaluate the contribution of MaxInfo to video understanding, we employed widely-used video understanding benchmarks, covering short-video tasks (such as MVBench [mvbench]) and medium-to-long video tasks (such as EgoSchema [egoschema], Video-MME [video-mme], and LongVideoBench [longvu]). For complete fairness, we only compared improved versions of the model against itself without MaxVol with freezing generation parameters, seed and prompt.
@@ 4.1 Main Results
Overall Performance. Table 1 present the performance gains achieved by integrating MaxInfo into existing InternVL2 [internvl], Qwen2-VL [qwen2-vl] and LLaVA-Video [llava-video] models. Experimental results show that MaxInfo Block exhibits significant performance improvements on a number of models. In particular, in the LLaVA-Video-7B and Qwen-VL-2B models, the introduction of MaxInfo Block improves the accuracy by 0.9%/1.7%, 6.4%, 3.3% and 1.4%/1.2%, 2.3%, 1.5% in VideoMME [video-mme], EgoSchema [egoschema], and LongVideoBench [longvideobench], respectively, which is significantly better than the versions without the MaxInfo block. This improvement not only validates the effectiveness of MaxInfo Block, but also provides new ideas for future research on video understanding tasks.
Table 1: Comparison of VLLM with and without MaxInfo on multiple benchmarks, where wo. sub. is without subtitles and with w. sub. subtitles.
@@
LLaVA-Video [llava-video] 7B 63.3/69.7(64) 57.3(64) 58.2(64)
+ MaxInfo 7B 64.2/71.464→(6,64) 63.7128→(12,64) 61.5128→(1,64)
[…]
LLaVA-Video [llava-video] 72B 70.5/76.9(64) 65.6(64) 61.9(64)
+ MaxInfo 72B 70.9/77.664→(6,64) 69.4128→(12,64) 64.9128→(1,64)
[…]
Qwen2-VL [qwen2-vl] 2B 55.6/60.4(786) 54.9(180) 47.3(256)
+ MaxInfo 2B 57.0/61.6256→(4,254) 57.2180→(12,180) 48.8256→(1,224)
[…]
Qwen2-VL [qwen2-vl] 7B 63.3/69.0(768) 66.7(180) 53.7(256)
+ MaxInfo 7B 62.1/70.0256→(4,254) 64.3180→(12,180) 55.7256→(1,224)
[…]
InternVL2 [internvl] 1B 43.0(16) 34.0(128) 43.4(128)
+ MaxInfo 1B 43.6128→(1,16) 34.1128→(1,32) 43.9128→(1,32)
[…]
InternVL2 [internvl] 2B 45.4(16) 46.1(128) 47.0(128)
+ MaxInfo 2B 45.0128→(1,16) 46.3128→(1,32) 47.3128→(1,32)
[…]
Although our results are slightly lower than the baseline on some models, we use significantly fewer frames than the baseline configuration.
Chunck Based MaxVol. Table 2 evaluates MaxInfo and Chunk-Based Scene-Awareness on the MLVU benchmark, showing consistent gains over uniform sampling. This approach is simple and entirely training-free, highlighting the potential of even basic scene-awareness. These results suggest that more advanced scene segmentation techniques could yield further improvements, making scene-awareness a promising direction for video understanding.
Table 2: Performance comparison on the MLVU benchmark between original Qwen2-VL and its variants with MaxVol and Chunk-Based. Best results in bold, second-best underlined.
@@
Comparison Baseline. As shown in Table 3, our proposed MaxInfo achieves strong competitiveness among training-free keyframe extraction methods and further demonstrates comparable or superior performance relative to existing state-of-the-art approaches.
Here, * indicates that our method does not require a predefined number of frames, but dynamically selects key frames based on the model’s initial inference frame count and the video’s information content. For example, on Video-MME and LongVideoBench, Qwen-MaxInfo processes an average of 180 and 170 frames, respectively, compared to 768 and 256 frames in the base Qwen model. Similarly, LLaVA-Video-MaxInfo uses 64 frames on both datasets, while the base LLaVA-Video model uses 58 and 51 frames, respectively.
Table 3: This table compares the performance of VLM with current state-of-the-art (SOTA) open-source and proprietary models, as well as key-frame selection methods, on video benchmark key-frame selection tasks. First place: bold, second place: underline, third place: italic.
Model LLM size #Frames VideoMME (wo sub.) EgoSchema LongVideoBench (val.)
Short Medium Long Overall
1.3min 9min 41min 17min 3min 12min
Proprietary Models
GPT4-o [openai2024gpt4o] - 1fps 77.1 62.1 59.2 66.2 - 66.7
Gemini-1.5-Pro [team2024gemini] - 1fps 82.3 75.3 67.5 75.7 - 64.0
Open Source Models
VideoChat2 [mvbench] 7B 16 48.3 37.0 33.2 39.5 - -
ShareGPT4Video [chen2024sharegpt4video] 8B 16 - - 37.9 43.6 - -
VideoLLaMA2 [cheng2024videollama] 7B 32 56.0 45.4 42.1 47.9 - -
LongVILA [xue2024longvila] 8B 128 60.2 48.2 38.8 49.2 - -
8B 256 61.8 49.7 39.7 50.5 - -
Qwen2-VL [qwen2] 7B 8 65.0 50.7 45.3 53.7 53.5 -
Key-frames Selection Methods
LongVU [longvu] 7B 1fps 64.7 58.2 59.5 60.6 67.6 -
LongVA [zhang2024long] 7B 16 59.0 46.6 43.6 49.7 - -
64 61.4 50.9 45.0 52.4 - -
128 61.1 50.4 46.2 52.6 - -
384 60.3 48.9 46.1 51.8 - -
Chat-Univi-v1.5 [Chat-univi] 7B 64 51.2 44.6 41.8 45.9 - -
AKS (LLaVA-Video) [tang2025adaptive] 7B 64 - - - 65.3 - 62.7
M-LLM (Qwen2-VL) [hu2025m] 7B - 69.6 54.1 51.9 58.7 65.9 -
BOLT (LLaVA-OneVision) [liu2025bolt] 7B 32 70.1 60.0 49.6 59.9 64.0 59.6
AdaReTaKe (Qwen2-VL) [wang2025adaretake] 7B - - - 56.4 64.2 - 57.2
AdaReTaKe (LLaVA-Video) [wang2025adaretake] 7B - - - 53.9 64.0 - 59.6
GenS (Qwen2-VL [yao2025generative]) 7B 50 - - - - - 58.7
ViLaMP [cheng2025scaling] 7B 1fps - - 58.4 67.7 - 60.2
Frame-Voyager [yu2024frame] 8B 128 (16) 67.3 56.3 48.9 57.5 - -
VideoTree [videotree] GPT-4 avg. 62 - - 54.2 - 61.1 -
Q-Frame (GPT-4o) [zhang2025q] GPT-4o 8 63.8 69.9 63.8 57.6 - 58.6
VSLS (GPT-4o) [guo2025logic] GPT-4o 32 71.9 61.9 55.2 63.0 - 63.4
VSLS (InternVL2.5-78B) [guo2025logic] 78B 8 59.0 57.5 57.7 58.1 - 64.5
MaxInfo (ours)
MaxInfo (InternVL2) 1B * - - - - - 43.9
MaxInfo (Qwen2-VL) 7B * 72.5 62.0 51.8 62.1 64.3 55.7
MaxInfo (LLaVA-Video) 7B * 74.6 63.3 54.6 64.2 63.7 61.5
MaxInfo (LLaVA-Video) 72B * 80.2 67.7 62.7 70.2 69.4 64.9